How many combination of three digits can be made from the number 0-9 that's larger than 15?
We notice that you said "combinations of three digits". You didn't say "3-digit numbers".
There are (10 x 10 x 10) = 1000 combinations of three digits.
The only combinations of 3 digits that form numbers less than 15 are 000, 001, 002, . . .
012, 013, and 014. There are 15 of those.
So the remaining (1000 - 15) = 985 are all larger than 15.
Your question does not have enough information to answer.
You cannot get an answer with a mathematical formula because you haven't added nearly enough information. Perhaps you could write a legal brief using an imaginary case and then the contributing lawyers could read it and see if any statistics would fit in anywhere, if they have the time. All you can get is a general answer that statistics often are used in litigation but they need to fit the facts of the case and be helpful to the plaintiff's or defendant's arguments. The attorney would more likely use a well authored statistic rather than create a mathematical formula.
Your question does not have enough information to answer.
You cannot get an answer with a mathematical formula because you haven't added nearly enough information. Perhaps you could write a legal brief using an imaginary case and then the contributing lawyers could read it and see if any statistics would fit in anywhere, if they have the time. All you can get is a general answer that statistics often are used in litigation but they need to fit the facts of the case and be helpful to the plaintiff's or defendant's arguments. The attorney would more likely use a well authored statistic rather than create a mathematical formula.
Your question does not have enough information to answer.
You cannot get an answer with a mathematical formula because you haven't added nearly enough information. Perhaps you could write a legal brief using an imaginary case and then the contributing lawyers could read it and see if any statistics would fit in anywhere, if they have the time. All you can get is a general answer that statistics often are used in litigation but they need to fit the facts of the case and be helpful to the plaintiff's or defendant's arguments. The attorney would more likely use a well authored statistic rather than create a mathematical formula.
Your question does not have enough information to answer.
You cannot get an answer with a mathematical formula because you haven't added nearly enough information. Perhaps you could write a legal brief using an imaginary case and then the contributing lawyers could read it and see if any statistics would fit in anywhere, if they have the time. All you can get is a general answer that statistics often are used in litigation but they need to fit the facts of the case and be helpful to the plaintiff's or defendant's arguments. The attorney would more likely use a well authored statistic rather than create a mathematical formula.
What does all you can eat data mean?
Well think about it this way: If you were having an all you can eat buffet, you'd eat as much as you could until you had no room left. It's pretty much the same, there's a limit, but one which is only reachable if you overuse the opportunity.
What is the largest 3 digit number made from 1 2 3 4 5?
543
well not exactly
it depends if u r allowed to use a number more than once it would probs b 555 if ur allowed to use it more than once
What is the definition of mean of all values in a population?
the mean is the average of the values of a population. the mean of 2, 8, 11 would be 7 while the median (equal numbers of values above and below) would be 8
What does median mean in kid form?
To find a median, put the set of numbers in least to greatest. The number in the middle = median.
What type of charts can be used to express data?
you can use a pie chart, a graph chart, or a graph chart. Hope i helped!
What is the percentage of not getting the sum of 3 on two dice?
Of the 36 possible outcomes of a throw of two dice, all but two result in a sum other than 3.
So Prob(Not a three) = 100*34/36 = 94.44... %
If you Roll a 4 or higher on a six-sided number cube what is the probablity?
The probability of rolling a 4 or higher on a standard die is 1 in 2.
I could not find a good derivation of the mgf for the normal on the web. However, I found a complete derivation on page 270 of Probability and Statistics by DeGroot, 2002 Edition. Most university libraries should have a copy.
The first equation on page 270 goes from exp(a)exp(b) to exp(a+b).
The second equation is a direct result of "completing the squares" method. See page 281.
In third equation on page 270 DeGroot goes from exp(a+b) to exp(a)exp(b) which allows to define a constant exp(mu*t +1/2*sigma^2 * t^2) which he can pull out from the integral.
The "disappearing" of the square-root with 2pi occurs because the C as defined in DeGroot iin the fourth equation is the complete integral of the cdf of a normal distribution, for any value of t.
DeGroot states we can replace mu with mu+sigma^2*t, but I would have preferred
Let mu* = mu +sigma^2*t, then for any t, the cdf of the normal, N(mu*, sigma), integrated from minus infinity to positive infinity must equal 1.
Thus what appeared to be an integration requires no integration at all. I could restate the derivation as:
mgf(t) = f(t) * integral of N(x, mu*,sigma) from minus infinity to positive.
mgf(t) = f(t) * 1
mgf(t) = exp(mu*t +1/2*sigma^2 * t^2)
Sometimes the mean does not give the entire picture of how the data behaves. For example: {2, 2, 3, 15, 100}
The mean is 24, the median is 3. This shows that if you graph these numbers on a line, they will be grouped much more to the low end than near the mean. ■
You use the sample mean as the estimate of the population mean; then as standard practice place confidence limits on the mean; most use 95%.
What are the possible numbers if the value of the tens digit is 60?
There are actually an infinite number of them. But if we leave out decimals
and negative numbers and stick exclusively to whole numbers, then there
are only ten. They are ... 60
61
62
63
64
65
66
67
68 and
69 . Oh no ! Wait a minute. Hold up, hee haw, whoa ! I'm wrong. Even if you stick to
only whole numbers, there are still an infinite number of them. Here are a few more
examples. The value of the tens digit is 60 in all of these: 160
161
263
462
1,065
9,760
21,869
314,167
2,718,281,828,459,065
.
.
etc.
If somebody rolls 2 standard six-sided dice what is the probability he obtains a sum of 7 or 11?
2 out of 12
Ans:
1 2 3 4 5 6 1 2 3 4 5 6
total no N of sums of two numbers assuming dice are indistinguishable:
N = 36
ways to get 7:
1 and 6 with a chance of (1/6)*(1/6) = 1/36
2 and 5 wih a chance of 1/36
4 and 3 with a chance of 1/36
whats the chance of getting 7 with either of these combinations:
1/36 + 1/36+ 1/36 = 1/12
what about 11:
6 and 5, which is the only way to get an 11:
the prob of this is (1/6)*(1/6) = 1/36
the prob of getting a 7 or an 11, with two dice:
1/36 + 1/12 = 1/9!
Hopefully someone will jump in if I've made a mistake.
OK, I've been told the answer is 2/9. If you do the same working as above, but you'll have to assume there are two ways to get each sum:
example, if you get a 7 from 2 and 5, then you can also get 7 from a 5 and 2. This means the dice are distinguishable!
If the specifications for a certain part is 8 decimeters long what is the length in meters?
0.8 metres (80 centimetres)
Tip: 1 decimetre is 10 centimetres